A 3 kg block moving to the right on a frictionless table at 4 m/s makes a head on collision with another 3 kg block moving 5 m/s to the left.


1
Expert's answer
2022-04-14T11:18:01-0400

The law of conservation of momentum says

"m_1v_1-m_2v_2=(m_1+m_2)u"

The speed of blocks after collision

"u=\frac{m_1v_1-m_2v_2}{m_1+m_2}=\frac{3*4-3*5}{3+3}=-0.5\:\rm m/s"

The velocity is directed leftward.